An oil exploration firm finds that 5% of the test wells it drills yield deposit of natural gas. If it drills 6 wells, find the probability that at least one well will yield gas.
Answer by greenestamps(13216)



P(at least one of 6) = 1 - P(none of 6)

P(none of 6) = (1-0.05)^6 = (0.95)^6

ANSWER: 1-(0.95)^6

Use a calculator....